15 books are being bought. Some cost $8.50 each, the remainder cost $9.50 each. Estimate the least number of books that can be bought at $8.50 each if the total cost must not be more than $135.50.
Answer by Nate(3500)

x = number of cheaper books
y = number of more expensive books
x + y = 15 or y = 15 - x
8.5x + 9.5y =< 135.5
Plug:
8.5x + 9.5(15 - x) =< 135.5
8.5x + 9.5(15 - x) =< 135.5
8.5x + 142.5 - 9.5x =< 135.5
-x =< -7
x >= 7
Atleast 7 books ... least amount being seven